I have a 61. If the final is 15% of my grade, what do I need to make to get a 80?
Answer by Alan3354(69443)

I have a 61. If the final is 15% of my grade, what do I need to make to get a 80?
=====================
61*85 + 15*F = 100*80
========================
If you make 100 on the final:
61*85 + 15*100 = 100*Grade
Grade = 66.85%
Good luck with that.